# RoboSpatial: Teaching Spatial Understanding to 2D and 3D Vision-Language Models for Robotics

### v1

@@ 3.2.2 Question-Answer Generation
Based on the generated relationships, we construct three types of question-answer pairs corresponding to each spatial relationship type: spatial configuration, spatial context, and spatial compatibility. To ensure that the model reasons with vision rather than relying on language commonsense, we structure the questions and answers to be template-based and unambiguous. We use a template: {object/space} {relationship} {anchor object} {reference frame}, where the spatial relationship and reference frame is defined in  Sec. 3.2.1. We maintain a balanced dataset with respect to spatial relationships to avoid introducing bias into the model. Furthermore, we use the 2D bounding box to create an auxiliary object-referring dataset to teach models what the object is, in order to avoid errors arising from a lack of object grounding. Using our pipeline we generated 3M spatial relationships, which is an order of magnitude larger than other spatial reasoning datasets shown in Table 1. Figure 2 shows an overview of the generated question types.
@@ 4 Experiments
4.1 Implementation
We apply our data generation process to three scene datasets and two tabletop datasets. The scene datasets include ScanNet [11], Matterport3D [7], and 3RScan [52]. The object tabletop datasets are HOPE [50] and GraspNet-1B [15]. We retrieve the 3D bounding box annotations and embodied images from EmbodiedScan [54]. This mix of datasets provides coverage for diverse indoor scenes featuring large objects for navigation, and smaller objects for learning priors for robot manipulation. In total, we generate a large scale dataset for spatial reasoning that contains around 3M spatial QA pairs with 5K 3D scans and 1M images. Statistics of these datasets are described in Table 3.
@@ 4.1 Implementation
Indoor Train 4914 scans 880K images 2.6M
Validation 50 scans 1K images 3K
Test 50 scans 1K images 3K
Tabletop Train 150 scenes 75K images 217K
Validation 40 scenes 3K images 3K
Test 77 scenes 3K images 3K
@@
Figure 3: In-domain and out-of-domain results for RoboSpatial-trained models. Two models are shown: SL (SpaceLLaVA [8]) and RP (RoboPoint [57]), where the S- prefix and FT indicates RoboSpatial-trained models. For Yes/No questions, the green checkmark indicates the correct answer. For spatial context questions, GT indicates the correct answer. For out-of-domain, all images except for the top right are from RoboSpatial-Home.
@@ 4.1.1 Trained 2D/3D VLMs
2D VLMs.  For our dataset, we want to compare a suite of VLMs, including ones fine-tuned for spatial understanding and ones that are not. We chose two base VLMs: VILA-1.5-8B [30] and LLaVA-NeXT-8B [34]; three specialized VLMs: SpaceLLaVA-13B, a community implementation of SpatialVLM [8]; RoboPoint-13B [57], a model fine-tuned on synthetic datasets to predict points in empty areas referenced by an object; and Molmo-7B [12], a new family of VLMs specializing in pointing and counting We also include a closed-source model, GPT-4o [40]. We omit models such as SpatialRGPT [9] due to its reliance on external object mask generation as input, which omits the important aspect of object grounding.
3D VLMs. Extending VLMs to the 3D domain introduces unique challenges due to the inherent complexity of 3D data. The scarcity of large-scale, high-quality 3D datasets hinders the development of end-to-end 3D VLMs. As such, there are few open source VLMs for 3D understanding that reason from 3D inputs. Among the available models, we selected: 3D-LLM [18], which lifts multi-view images into a 3D point cloud, and LEO [20], which directly takes as input 3D object point clouds by relying on a pre-processing object segmentation step. These models allow us to evaluate spatial understanding in 3D by processing data either through reconstructed 3D representations or directly from point clouds. By incorporating both approaches, we aim to assess the effectiveness of 3D VLMs in handling spatial information from different types of 3D data inputs.
Fine-tuning. Our evaluation includes both “zero-shot” evaluation of each VLM, plus evaluation after fine-tuning on RoboSpatial for open-source VLMs. See the Appendix for further details.
4.1.2 Spatial Understanding Evaluation
We evaluate our method in answering 3000 questions (1000 questions per spatial relationship) from our test set, and measure how often the model produces the correct answer. The evaluation includes binary yes/no questions, and questions requiring numeric responses. For yes/no questions, we report the success rate in producing the correct response. For numeric questions, we report the rate at which the produced numeric answer is within the convex hull of a set of points generated by our spatial relation extraction method. Note that this is potentially an overly strict metric, and therefore the reported performance on numeric questions reflects a lower-bound of the true performance. Table 4 shows the main results.
4.1.3 Cross-Dataset Generalization Evaluation
Next, we evaluate the generalization capability of our method by testing it across different scene types—specifically, both indoor and tabletop scenes—to control for any bias in the annotations of the underlying datasets that make up our benchmark. We train on data derived from subsets of the datasets corresponding to one scene type (either indoor or tabletop) and test on held-out datasets from the other scene type, representing unseen environments. We expect that even when training on a subset of datasets, the performance on unseen scene types will improve if our method generalizes well. The results of this cross-dataset evaluation are shown in Table 5.
4.1.4 Out-of-Domain Evaluation
To test the out-of-domain transferability of RoboSpatial-trained models, we evaluate them on RoboSpatial-Home, our manually annotated indoor scene dataset, and the validation spatial reasoning split of the Blink [17] benchmark.
RoboSpatial-Home is a collection of 50 crowd-sourced indoor RGBD images captured using a handheld iPhone camera equipped with a built-in depth sensor. Each image is manually annotated with three spatial relationships, resulting in a total of 150 questions across 50 images. We designed RoboSpatial-Home to test the RoboSpatial-trained model’s capability to generalize to unseen objects in real life cluttered scenes.
Blink [17] is a visual reasoning benchmark designed to evaluate the reasoning capabilities of vision-language models (VLMs) using visual input alone. Blink was chosen to assess the generalization ability of RoboSpatial-trained models on unseen spatial relationships such as “next to,” “touching,” and “on top.” Unlike benchmarks such as What’s Up [26], which include spatial relationships similar to those in RoboSpatial, Blink [17] provides a more rigorous test of out-of-domain generalization. The validation spatial reasoning split of Blink consists of 143 binary yes/no questions about spatial configurations.
The results of the out-of-domain generalization evaluations are presented in Table 6 and visualized in Fig. 3.
@@ 4.1.4 Out-of-Domain Evaluation
Table 4: Results of existing 2D/3D VLMs on a held-out test split of images and scans. All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on our RoboSpatial dataset. The best result for each column is bolded.
Indoor $\rightarrow$ Tabletop Tabletop $\rightarrow$ Indoor
@@
[…]
Model RoboSpatial-Home BLINK-Spatial
Localization Affordance Compatibility Accuracy
Open-source VLMs
2D VLMs
VILA [30] 53.3 12.0 52.0 72.7
+RoboSpatial 62.0 $\uparrow$ 32.0 $\uparrow$ 58.0 $\uparrow$ 79.7 $\uparrow$
LLaVA-Next [34] 48.0 9.3 37.3 71.3
+RoboSpatial 58.0 $\uparrow$ 24.0 $\uparrow$ 44.0 $\uparrow$ 79.0 $\uparrow$
SpaceLLaVA [8] 60.0 16.0 49.3 76.2
+RoboSpatial 68.7 $\uparrow$ 38.0 $\uparrow$ 56.0 $\uparrow$ 81.8 $\uparrow$
RoboPoint [57] 43.3 41.3 36.0 63.6
+RoboSpatial 50.0 $\uparrow$ 54.0 $\uparrow$ 48.0 $\uparrow$ 70.6 $\uparrow$
3D VLMs
3D-LLM [18] 40.0 8.0 46.0 N/A
+RoboSpatial 48.0 $\uparrow$ 36.0 $\uparrow$ 52.7 $\uparrow$ N/A
LEO [20] 50.7 10.0 48.0 N/A
+RoboSpatial 64.0 $\uparrow$ 40.0 $\uparrow$ 60.0 $\uparrow$ N/A
Not available for fine-tuning
Molmo [12] 44.7 38.0 58.0 67.1
GPT-4o [40] 64.0 25.3 56.7 76.2
4.2 Results
Our experiments on the held-out test split in Table 4, out-of-domain splits in Table 6, and real-world robot experiments in Table 7 clearly demonstrate the power of RoboSpatial in teaching spatial reasoning to VLMs where training on RoboSpatial improved spatial understanding across all spatial reasoning benchmarks. We further answer the following detailed questions about  RoboSpatial-trained models:
How well does RoboSpatial training generalize to unseen spatial relationships? While RoboSpatial contains templated question-answer pairs and a fixed set of spatial relationship prepositions, Table 6 shows that RoboSpatial training does generalize to other spatial relationship prepositions present in Blink [17], such as “under,” “next to,” and “far away.” We highlight that RoboSpatial encompasses all six principal directions along the x, y, and z axes in 3D space, and that generalizing to novel relationships involves mapping each of these directions to one or more spatial relationship prepositions, a task at which LLMs excel. For example, “on top of” and “on” refer to the “above” direction in a world-centric frame, while “under” corresponds to “below.” Furthermore, we observe that prepositions such as “next to” and “beside” imply that objects need to be in close proximity to one another. Since RoboSpatial contains spatial context questions that teach the model to generate points close to a given object, this implicitly teaches what it means to be “close” to another object, enabling RoboSpatial-trained models to understand concepts like “next to.” However, RoboSpatial training does not seem to generalize well to questions involving human-centric perspectives, since we focus solely on objects in our dataset for robotics use cases.
Do RoboSpatial-trained models understand nuanced perspectives? When humans refer to spatial relationships, we often imply a specific frame of reference. For example, “in front of the car” typically refers to the car’s front hood. Therefore, questions in RoboSpatial-Home omit explicitly specifying the reference frame to test the model’s ability to align its understanding with the questioner’s intended reference frame. Results in Table 6 show that models trained with RoboSpatial can automatically infer the reference frame, with Figure 3 showing examples like “Is the frame in front of the nightstand.”
Are 3D VLMs better at learning spatial relationships than 2D VLMs? While our results seem to imply that 3D VLMs outperform their 2D counterparts in learning spatial relationships, we point out that 3D-LLM [18] and LEO [20] are trained with indoor scan datasets, making the comparison between 2D and 3D models not entirely fair. Nevertheless, we hypothesize that 3D models have an advantage in spatial understanding due to their ability to incorporate depth information. Therefore, we developed RoboSpatial to be both 2D and 3D ready to support further research in this area.
Model Success Rate (%)
@@ 4.2 Results
Table 7: Success rate comparison between robotics systems powered by zero-shot VLMs and VLMs fine-tuned on RoboSpatial.
@@ 4.3 Real Robot Experiments
We designed a comprehensive suite of tasks that require spatial reasoning and object manipulation. We use a diverse set of objects, including various shapes (e.g., cubes, cylinders, boxes), colors (red, orange, blue, green), and everyday items (food items, toys) as shown in Figure 4. Our goal in selecting these objects is not to confound the methods with difficult-to-identify items but to make identification as simple as possible with no ambiguity. For each scene configuration, we asked two yes/no questions, followed by asking a pick-and-place question. We note that picking and placing were treated as independent tasks for evaluations. We conducted over 200 model queries during our experiments, where details of questions and scenes used in the experiments are in the Appendix. For robot experiments, we used a Kinova Jaco robot [6] paired with a ZED2 camera for RGB(D) perception and employ cuRobo [47] for collision-free motion planner.
We evaluate the following VLMs: LLaVA-NeXT [34] and RoboPoint [57], both with and without RoboSpatial training; and Molmo [12] and GPT-4o [40] as baselines. Table 7 and Fig 4 presents the results of the real-world robot experiments. Our results show that for generic VLM such as LLaVa-Next [34], training on RoboSpatial significantly improved its spatial understanding capabilities in robot manipulation tasks. Furthermore, RoboSpatial-trained models were more likely to align predictions with the implied reference frame, as demonstrated in the “place in front of the pony” task, where only RoboSpatial-trained models’ predictions aligned correctly in front of the pony’s head. Similarly, in the “place in front of the orange juice box” task, RoboSpatial-trained models could identify an appropriate distance to place the object relative to its size. In contrast, for RoboPoint, the predicted point was too far to the object, making it difficult to place the object correctly. However, we note that GPT-4o performs comparably with RoboSpatial-trained RoboPoint. We attribute this to GPT-4o being able to understand much more diverse language inputs, unlike RoboSpatial, which was trained on templated language.
Here, we show that our dataset enhances the performance of 2D VLMs in robotics tasks. In our robot setup, 2D VLMs project predicted points from pixel space into 3D space; however, even a 2-pixel error can translate into a displacement of 5 to 10 cm in 3D space, leading to task failure. Interesting future direction is exploring how perspective impacts 2D point selection in the context of VLMs. Another interesting future direction is to enable 3D VLMs to capably support partial 3D point cloud observations, unlike the models we currently train, which require complete colored 3D point clouds, so that they can be effectively utilized in real-world robotics scenarios.

### v2

@@ 3.2.3 Question-Answer Generation
Once the spatial relations $\{s_{i}\}$ have been extracted, we generate corresponding question-answer pairs $\{d_{i}\}$ using structured templates. Each question follows the format: {target} {relation} {anchor} {ref. frame} where the relation and frame are defined in Section 3.2.1. To ensure that models learn from visual grounding rather than linguistic priors, we use deterministic templates that avoid ambiguity and minimize reliance on commonsense.
Each spatial relation type—context, compatibility, and configuration—has a corresponding question format. Configuration and compatibility tasks result in binary (True/False) answers. Context questions produce a list of valid 2D coordinates in image space.
Correctly resolving which object is being referred to in a spatial question is essential for reliable spatial understanding. To reduce errors arising from incorrect object identification, we additionally generate an auxiliary object grounding dataset that links object descriptions to 2D bounding boxes in the image. These grounding annotations are derived by projecting 3D bounding boxes into image space using camera intrinsics and extrinsics. This supervision helps models more accurately resolve references during spatial reasoning and is included during training. See Appendix B.3 for details.
Using this pipeline, we generate around 3 million spatial relationships and their associated question-answer pairs. This scale is an order of magnitude larger than prior spatial reasoning datasets (see Table 1).
@@ 4 Experiments
4.1 Setup
We apply the data generation pipeline to three scene datasets—ScanNet [8], Matterport3D [4], and 3RScan [56]—and two tabletop datasets—HOPE [54] and GraspNet-1B [12]. We retrieve 3D bounding box annotations and embodied images from EmbodiedScan [58], and generate a large-scale spatial reasoning dataset covering diverse indoor environments: larger scenes for navigation and smaller object-centric setups for manipulation.
In total, RoboSpatial includes approximately 3M spatial QA pairs across 5k 3D scans and 1M images. (Table 2 provides a breakdown; values are rounded to the nearest thousand for clarity.)
@@ 4.1 Setup
Indoor Train 4916 scans 883k images 2.7M
Validation 40 scans 1k images 3k
Tabletop Train 190 scenes 76k images 220k
Validation 77 scenes 355 images 3k
@@ 4.1.1 Trained 2D/3D VLMs
2D VLMs.  We evaluate several vision-language models (VLMs) using RGB-only image inputs. Our selected base VLMs are VILA-1.5-8B [30] and LLaVA-NeXT-8B [35]. We also include three specialized models: SpaceLLaVA-13B (a community version of SpatialVLM [5]), RoboPoint-13B [62] (trained to predict points in empty space given an object reference), and Molmo-7B [9] (designed for pointing and counting). We also include GPT-4o [42] as a closed-source baseline. We omit models such as SpatialRGPT [6] that depend on external mask inputs, as they bypass the object grounding challenge.
3D VLMs.  Models operating over 3D data must handle richer, more complex spatial representations. We include two models that process 3D inputs: 3D-LLM [18], which reconstructs colored 3D point clouds from multi-view RGB images, and LEO [20], which operates on segmented colored point clouds of individual objects. These models allow us to explore spatial reasoning when models consume RGBD or point cloud representations directly.
Fine-tuning.  We evaluate models in both zero-shot and fine-tuned settings, using RoboSpatial to fine-tune open-source models. To mitigate failure cases arising from poor object grounding, we also include an auxiliary grounding dataset during training, which provides additional supervision for object reference resolution. This auxiliary dataset does not contribute to spatial reasoning performance. (See Appendix for ablation experiments.)
@@
Table 3: Results of existing 2D/3D VLMs on a held-out validation split (RoboSpatial-Val) of images and scans. All methods, for all tasks, perform better ($\uparrow$) when fine-tuned on RoboSpatial. The best result for each column is bolded.
Model RoboSpatial-Home BLINK SpatialBench
Configuration Context Compatibility Accuracy Accuracy
2D VLMs
VILA [30] 57.8 0.0 69.0 72.7 53.0
+RoboSpatial 65.9 $\uparrow$ 15.6 $\uparrow$ 78.0 $\uparrow$ 79.7 $\uparrow$ 73.6 $\uparrow$
LLaVA-NeXT [35] 60.2 0.0 71.4 71.3 55.9
+RoboSpatial 76.4 $\uparrow$ 19.7 $\uparrow$ 80.1 $\uparrow$ 79.0 $\uparrow$ 70.6 $\uparrow$
SpaceLLaVA [5] 61.0 0.08 34.3 76.2 47.1
+RoboSpatial 71.6 $\uparrow$ 13.1 $\uparrow$ 62.0 $\uparrow$ 81.8 $\uparrow$ 67.7 $\uparrow$
RoboPoint [62] 59.4 2.5 80.1 63.6 44.1
+RoboSpatial 70.0 $\uparrow$ 21.3 $\uparrow$ 88.6 $\uparrow$ 70.6 $\uparrow$ 64.7 $\uparrow$
3D VLMs
3D-LLM [18] 39.8 0.0 35.2 N/A N/A
+RoboSpatial 55.2 $\uparrow$ 8.2 $\uparrow$ 52.3 $\uparrow$ N/A N/A
LEO [20] 51.2 0.0 38.1 N/A N/A
+RoboSpatial 64.2 $\uparrow$ 10.0 $\uparrow$ 57.1 $\uparrow$ N/A N/A
Not available for fine-tuning
Molmo [9] 48.7 0.0 24.8 67.1 55.9
GPT-4o [42] 74.0 6.6 55.2 76.2 70.6
4.1.2 Spatial Understanding Evaluation
We evaluate spatial reasoning capabilities using RoboSpatial-Val, a held-out validation subset of RoboSpatial sampled from scans that are entirely unseen during training. This benchmark comprises 6,000 heuristically generated questions from our data generation pipeline, with 2,000 questions per spatial relation type. Questions fall into two categories: binary yes/no questions and coordinate prediction tasks. For yes/no questions, we report accuracy. For coordinate predictions, we evaluate whether the model’s predicted 3D location lies within the convex hull of a reference point set derived from scene geometry.
While this convex hull criterion provides a well-defined geometric target, it is arguably overly strict—e.g., predictions near but just outside the boundary are marked incorrect. As a result, reported scores represent a conservative estimate of each model’s spatial understanding. RoboSpatial-Val serves as the primary benchmark for comparing 2D and 3D VLMs trained on RoboSpatial, enabling controlled evaluation within the same data distribution. Results are presented in Table 3.
4.1.3 Cross-Dataset Generalization Evaluation
To assess generalization across environment types, we partition the training data into indoor-scene and tabletop subsets. Models are trained on one type and evaluated on held-out datasets from the other. Despite differing object distributions and scene layouts, we observe a positive synergy between indoor and tabletop environments: training on one environment type improves spatial reasoning on the other, as shown in Table 6.
Figure 3: In-domain (RoboSpatial-Val, top) and out-of-domain (RoboSpatial-Home, BLINK [15], middle and bottom) results for RoboSpatial-trained models. Two models shown: SL (SpaceLLaVA [5]) and RP (RoboPoint [62]); the -FT suffix indicates fine-tuning on RoboSpatial. Correct answers in green. All images except bottom-right in the out-of-domain rows are from RoboSpatial-Home.
4.1.4 Out-of-Domain Evaluation
To test the out-of-domain transferability of RoboSpatial-trained models, we evaluate on three benchmarks: RoboSpatial-Home, BLINK [15], and SpatialBench [2]. RoboSpatial-Home contains 350 manually written spatial questions over diverse real-world RGBD scenes captured with an iPhone equipped with a depth sensor. We curated this benchmark to evaluate generalization to novel indoor settings with previously unseen objects. BLINK is a visual reasoning benchmark consisting of binary spatial questions involving relationships such as “next to,” “touching,” and “on top.” BLINK allows us to assess the ability of models to assess generalization of spatial reasoning to unseen language configurations. We evaluate only the spatial portion of BLINK, as that aligns with the core focus of RoboSpatial. SpatialBench, introduced in the SpatialBot [2] paper, uses RGB inputs and tests spatial understanding across several categories. We focus on the position category, which most directly aligns with our emphasis on spatial localization and placement. Together, these benchmarks complement RoboSpatial-Val by covering a broader range of visual and linguistic variation.
4.2 Results
We evaluate the effectiveness of RoboSpatial in improving spatial reasoning capabilities in VLMs across held-out and out-of-domain benchmarks. In this section, we focus on analyzing the model’s generalization and understanding of spatial relationships. We address the following questions:
How well does RoboSpatial training generalize to unseen spatial relationships? Although RoboSpatial consists of template-generated QA pairs with a fixed set of spatial prepositions, we observe in Tab. 4 that models trained on it can generalize to spatial relationships not explicitly included in the training set. This is particularly evident in evaluations on the BLINK dataset [15], which contains diverse prepositions such as “under,” “next to,” and “far away.” We attribute this generalization to the fact that RoboSpatial encompasses all six principal directions in 3D space (along the x, y, and z axes). Generalizing to new prepositions often requires mapping linguistic expressions (e.g., “on top of,” “under”) to these spatial primitives—a task at which LLMs are naturally proficient. For example, “on top of” often refer to “above” in a world-centric frame, while “under” maps to “below.” Moreover, prepositions such as “next to” or “beside” imply proximity between objects. Because RoboSpatial includes questions that require generating points near a reference object, it implicitly teaches the concept of closeness. This enables trained models to understand these proximity-based relationships, even if they are not explicitly represented during training.
Do RoboSpatial-trained models understand nuanced perspectives? Spatial references in natural language often imply specific reference frames. For instance, “in front of the car” typically refers to the direction of the car’s front hood. In RoboSpatial-Home, we omit explicit frame specifications in the questions to evaluate whether models can align with the implicit reference frame intended by the questioner. We find that models trained with RoboSpatial can often infer the correct frame of reference, suggesting that they have learned to associate object geometries and orientations with spatial language. Figure 3 shows examples such as “Is the frame in front of the window?”, where the model accurately identifies the intended spatial relation.
Are 3D VLMs better at learning spatial relationships than 2D VLMs? The findings in Tab. 3 suggest that 3D VLMs tend to outperform 2D counterparts in spatial reasoning tasks, likely due to their ability to directly utilize depth information. However, this comparison is not entirely fair: models like 3D-LLM [18] and LEO [20] are pretrained on RGB-D indoor scan datasets, some of which overlap with the environments used in the source datasets (e.g., Matterport3D, ScanNet). This gives them prior exposure to scene geometry and object layouts, which may bias their performance. To support more controlled and fair comparisons in the future, we designed RoboSpatial to be compatible with both 2D and 3D modalities, allowing researchers to investigate the impact of modality, architecture, and pretraining data under unified evaluation protocols.
Indoor Tabletop
$\downarrow$ $\downarrow$
Tabletop Indoor
@@ 4.2 Results
Table 5: Cross-dataset generalization results between indoor and tabletop environments.
Model Success (%)
@@
Table 6: Robot experiment results.
@@ 4.3 Real Robot Experiments
We design a suite of tabletop manipulation tasks requiring spatial reasoning. The setup includes a Kinova Jaco robot [3], paired with a ZED2 camera for RGB-D perception. The robot system implements actions to pick and place objects on the table using cuRobo [52] for motion planning. Tasks include spatial questions that require a yes/no answer, and pick-and-place instructions that require successfully controlling the robot to complete the task. We adopt a modular design, where the VLM is queried for spatial understanding, and the resulting predictions (e.g., target points) are passed to a separate motion planning system for execution. We use a range of simple, unambiguous objects—colored cubes, cylinders, food items, and toys—to ensure the challenge lies in spatial understanding rather than object recognition (Figure 4). In total, we conducted over 200 model queries. Details of the questions and scene configurations are provided in the Appendix D.5. We evaluate the following VLMs: LLaVA-NeXT [35] and RoboPoint [62], both with and without RoboSpatial training; and two strong baselines, Molmo [9] and GPT-4o [42]. Table 6 and Figure 4 present the results.
Experiments show that LLaVA-NeXT fine-tuned on RoboSpatial achieves the highest success rate across all models. Training with RoboSpatial enhances spatial understanding in 2D VLMs, enabling the model to correctly interpret instructions such as “place in front of the pony,” where placement is aligned with the pony’s head direction. It also demonstrates sensitivity to object scale, as in the task “place in front of the orange juice box,” where the model places the object at a reasonable distance. In contrast, baseline models such as RoboPoint frequently place objects too far from the target, likely due to limited understanding of spatial proximity. We also observe that spatial failures in 2D VLMs often stem from errors in projecting 2D predictions into 3D. Even a small 2-pixel shift in image space can translate to a 5–10 cm error in the physical world, which is significant in manipulation tasks. Nonetheless, models trained on RoboSpatial produce more accurate predictions, reducing these failure cases and showing the benefit of dataset-driven improvements. Interestingly, GPT-4o performs comparably to RoboSpatial-trained RoboPoint. We attribute this to GPT-4o’s broader language understanding and instruction-following ability, which partially compensates for its lack of task-specific spatial training. Looking forward, promising directions include investigating how viewpoint affects 2D spatial predictions, and developing 3D VLMs that can reason over partial point clouds—removing the need for complete 3D scans and making deployment in real-world systems more feasible.